# Localization and splitting of a quantum droplet by immersing a heavy impurity

### v1

@@ Abstract
We unravel the existence and nonequilibrium response of one-dimensional harmonically trapped droplet configurations in the presence of a central repulsive or attractive potential well mimicking the effect of a heavy impurity. For fixed negative chemical potentials, it is shown that droplets fragment into two for increasing repulsive potential heights, a process that occurs faster for larger widths. However, atoms from the droplet accumulate at the attractive potential, especially for wider ones, leading to a deformed droplet and eventually to the termination of the solution. Linearization analysis yields the underlying excitation spectrum which dictates stability and the behavior of the ensuing collective modes. Quenches in the potential height are used to demonstrate dynamical fragmentation of the droplet for repulsive potentials as well as self-evaporation along with droplet localization and eventual relaxation for longer evolution times in the case of attractive potential wells. The many-body character of the dynamics is explicated by evaluating the participating single-particle eigenstates manifesting the superposition nature of the droplet state. Our results should be detectable by current ultracold atom experiments and may inspire engineered droplet dynamics with the aid of external potentials.
@@ I Introduction
Recent ultracold atom experiments offer access to the regime of strong attractive interactions Böttcher et al. (2020); Chomaz et al. (2023); Chen and Hung (2021). Such settings can host, among others, ultradilute liquid-type many-body droplet states Luo et al. (2021); Malomed (2021); Mistakidis et al. (2023) that are sustained due to the presence of quantum fluctuations in dipolar gases Böttcher et al. (2020); Politi et al. (2022); Chomaz et al. (2023) as well as homonuclear Cheiney et al. (2018); Semeghini et al. (2018); Cabrera et al. (2018) and heteronuclear D’Errico et al. (2019); Burchianti et al. (2020) short-range Bose mixtures. Commonly the interaction Böttcher et al. (2020) and dimension Ilg et al. (2018); Pelayo et al. (2025a) dependent Lee-Huang-Yang (LHY) first-order correction term Lee et al. (1957) to the mean-field energy suffices to capture the formation and dynamics of quantum droplets within a suitable extended Gross-Pitaevskii (eGPE) model Petrov (2015); Petrov and Astrakharchik (2016); Englezos et al. (2025). Beyond LHY effects have been, for instance, documented in Refs. Parisi et al. (2019); Mistakidis et al. (2021); Englezos et al. (2023).
@@
On the other hand, the immersion of impurity atoms into droplets can bring forth additional mechanisms including self-localization of the impurity Abdullaev and Galimzyanov (2020); Bighin et al. (2022), hybrid droplet-impurity modes Sinha et al. (2023), tunability of the dipolar bound state character Wenzel et al. (2018), and mediated impurity interactions from the droplet Pelayo et al. (2024). The impact of a heavy impurity on its host bosonic droplet remains, however, elusive. As demonstrated in repulsive Bose gases Campbell et al. (2014); Brauneis et al. (2022), the heavy impurity can be well approximated by an external potential barrier, an approach we also adopt in this work. In this context, a repulsive barrier dragged through the droplet has been used recently to dynamically generate defects like solitary waves and vortices Saqlain et al. (2023); dos Santos et al. (2025), a well-established protocol in repulsive Bose gases Engels and Atherton (2007); Hakim (1997); Rodrigues et al. (2009); Katsimiga et al. (2018) but with an amended local speed of sound. Additionally, the transmission and reflection properties of droplets passing through an attractive localized potential well have been reported Debnath et al. (2023).
@@
Accordingly, it is interesting to study the effect of the impurity onto the droplet such as to retain its bound state character in a structurally modified configuration or its destruction. Here, the computation of the underlying excitation spectrum being able to reveal the stability of the resulting states and their ensuing modes is certainly desirable. In another vein, the dynamical response of the droplet by switching-on the impurity can trigger specific excitation modes or even fragmentation. To explore these possibilities, we deploy an infinitely heavy impurity represented by a potential barrier which can be either attractive or repulsive, while being embedded into a symmetric Petrov and Astrakharchik (2016) (alias single-component) one-dimensional (1D) quantum droplet featuring contact interactions. Similar external confinements with attractive potentials are used to emulate the motion of electrons in solids Atkins and Friedman (2011), which further promotes the broader character of our results. The static properties and nonequlibrium dynamics of this setting are modeled by the appropriate eGPE Petrov and Astrakharchik (2016).
@@
On the ground-state, we determine regions of existence of droplet states in the presence of an attractive or a repulsive potential well. The former favors droplet’s spatial localization as was also argued in Ref. Debnath et al. (2023). However, here we explicate the controllability of this mechanism demonstrating, for instance, the existence of an upper bound of the height above which the droplet disappears. This threshold decreases for larger widths or greater LHY strengths, which is traced back to the appearance of additional bound states in the potential and the increased droplet localization Petrov (2015) respectively. In contrast, a repulsive potential well enforces the gradual deformation and eventual fragmentation of the droplet into two, a process that occurs faster for reduced heights in the case of larger widths. We also infer the spectral (in)stability of the ensuing droplet configurations for (repulsive) attractive potentials through Bogoliubov-de Gennes (BdG) linearization analysis. The behavior of their collective modes in terms of the potential well characteristics is also assessed.
@@
Upon quenching a repulsive potential well, we showcase droplet’s fragmentation into two counterpropagating segments whose separation and velocity can be tuned by means of the potential height and width. A semiclassical model is devised to capture this response, based on energy conservation. Meanwhile, the quench excites the droplet fragments which suffer self-evaporation Ferioli et al. (2020) releasing small density portions. On the other hand, the sudden ramp-up of the attractive well enforces spatial localization of the droplet which features excitation processes. The latter, due to self-evaporation, decay in the course of the evolution letting the droplet to eventually attain a quasi-steady state. Monitoring the position variance allows to track the collective motion of the droplet and its excitation frequencies whose values are in-sync with the BdG predictions. Finally, insights into the participating microscopic mechanisms and the crucial role of the quantum superposition for the droplet dynamics are unveiled through the single-particle eigenstates of the external potential. The dynamical behavior of the eigenstates inherits information about the droplet’s excitation processes.
@@
This work is organized as follows. Section II introduces our droplet setting along with its external potential and eigenstates as well as the reduced single-component eGPE framework. Section III elaborates on the characteristics of the droplet solutions and their stability in the presence of an attractive and a repulsive potential barrier. In Section IV we analyze the emergent nonequilibrium droplet dynamics induced by suddenly switching on the potential barrier which facilitates the dynamical splitting [Sec. IV.1] or localization [Sec. IV.2] of the original droplet. A decomposition of the many-body dynamics in terms of the single-particle eigenstates of the external confinement obtained through exact diagonalization is discussed in Sec. IV.3. We conclude and discuss future research extensions based on our results in Section V.
Figure 1: Schematic illustration of the external potential described by Eq. (2) corresponding to (a) a double-well with $V_{0}>0$ and (b) a harmonic trap with a central dip when $V_{0}<0$. The single-particle energy levels are marked by $\epsilon_{n}$, with $n=0,1,2,\dots$, obtained through exact diagonalization. The shaded areas in panel (a) designate the energy doublets. The respective eigenstates are provided in panels (c) [(d)] for $V_{0}>0$ [$V_{0}<0$]. In both cases, the employed potential characteristics correspond to $V_{0}=\pm 1$ and $\sigma=0.5$. (e), (f) The dependence of the underlying energy gaps among consecutive energy levels, $\Delta\epsilon_{n}=\epsilon_{n}-\epsilon_{n-1}$, with $n=1,2,\dots$ as a function of the potential depth $V_{0}$ (see legend). Inset in panel (d) shows a magnification at the center of $\left|\varphi_{4}\right|^{2}$ to clearly visualize the existent“kink” of even eigenstates. In the double-well, $\Delta\epsilon_{n}$ among doublets increase for larger $V_{0}$, while in the attractive potential case additional bound states (BS) appear for increasing $V_{0}$ (shaded regions).
@@ II Droplet setting and external potential
An elongated, across one spatial dimension, homonuclear ($m_{1}=m_{2}\equiv m$) bosonic mixture with intracomponent repulsion ($g_{11}=g_{22}\equiv g>0$) and intercomponent attraction ($g_{12}<0$) is assumed in order to form droplet configurations Petrov and Astrakharchik (2016). Droplets in 1D can be generated under the condition $\delta g=g_{12}+g>0$ Petrov and Astrakharchik (2016); Mistakidis et al. (2023). The participating interaction strengths are experimentally tunable through manipulation of the three-dimensional scattering lengths that can be achieved by means of available magnetic fields using Fano-Feshbach resonances Semeghini et al. (2018); Cheiney et al. (2018) or via confinement induced resonances Olshanii (1998); Bergeman et al. (2003) which can be attained by adjusting the transverse confinement. The two hyperfine states of this mixture are equally populated ($N_{1}=N_{2}\equiv N$) and hence they behave identically. Promising candidates for the realization of these components are the hyperfine states |$\ket{F=1,m_{F}=-1}$ and |$\ket{F=1,m_{F}=0}$ of 39K which have been used in relevant three-dimensional experiments Cabrera et al. (2018); Semeghini et al. (2018); Ferioli et al. (2019).
The mixture is confined by a harmonic oscillator with frequency $\omega_{x}\ll\omega_{\perp}$. This condition ensures the 1D character of the system, meaning that due to tight confinement along the perpendicular $y$, $z$ directions the atoms are kinematically constrained across the $x$ direction. Accordingly, the 1D droplet is adequately described by the eGPE Petrov and Astrakharchik (2016); Englezos et al. (2025)
@@
experiencing the competition of repulsive cubic mean-field interactions and attractive quantum fluctuations scaling as $\Psi^{2}$. The latter are responsible for droplet formation Petrov and Astrakharchik (2016); Luo et al. (2021).
The external potential consists of a loose harmonic trap [^1] Note that the presence of the harmonic trap is not necessary for the results to be presented below. Similar phenomena occur also with a box confinement in the $x$-direction. with $\Omega=\omega_{x}/\omega_{\perp}=0.001$ and a superimposed central barrier
@@
Here, $\sigma$ denotes the spatial width of the central potential, and $V_{0}$ refers to the strength of the potential well which can be attractive ($V_{0}<0$) or repulsive ($V_{0}>0$). Apparently, the central potential well reduces to a delta-like one for $\sigma\to 0$ and $V_{0}\to\pm\infty$, while $V_{0}\sigma$ is kept fixed. Specifically, for $V_{0}<0$, the overall confinement corresponds to a harmonic trap with a central dip, see Fig. 1(b), facilitating the accumulation of atoms from the droplet at its vicinity since it hosts additional bound states depending on $V_{0}$ (see also the discussion below). This contribution acts against flat-top configurations and its increase may gradually destruct the droplet. The potential hosts a larger number of bound states for larger $V_{0}$ [Fig. 1(f)], while their energy gap is reduced for smaller $\sigma$. On the other hand, for $V_{0}>0$ a 1D double-well potential [Fig. 1(a)] is realized which favors droplet fragmentation into two segments. In this case, a larger $V_{0}$ results in the increase of the energy gap among consecutive single-particle eigenstate doublets, see also Fig. 1(e).
@@
where $\lambda=\frac{\delta g}{2}\sqrt{(\frac{m}{\omega_{\perp}\hbar^{3}})}$, and $\delta=\frac{1}{\pi}(\frac{mg^{2}}{\omega_{\perp}\hbar^{3}})^{3/4}$ represent respectively the strength of the mean-field and the LHY quantum correction. Also, $V(x)=\frac{1}{2}\Omega^{2}x^{2}+\frac{V_{0}e^{-x^{2}/{2\sigma^{2}}}}{\hbar\omega_{\perp}}$. As such, all quantities to be presented below are in dimensionless units. Since we aim to examine the non-trivial impact of the potential well on the droplet, barrier heights satisfying $\left|V_{0}\right|\gg\left|\mu\right|$ (with $\mu$ denoting the droplet’s chemical potential) are considered. Also, the healing length of the droplet, $\xi=\sqrt{\lambda}/\delta=1$, sets a characteristic length scale. Accordingly, the potential widths examined obey $\sigma<\xi$; otherwise droplet destruction occurs already for very small $V_{0}$ values.
Typical trap ratios $\Omega=0.001$ can be achieved through tight transversal confinement $\omega_{\perp}/(2\pi)=1~{}$kHz which is in a similar range to the 1D experiments of Refs. Görlitz et al. (2001); Moritz et al. (2003); Meinert et al. (2017); Romero-Ros et al. (2024) and frequency $\omega_{x}/(2\pi)=1~{}$Hz along the elongated axis. The central potential well is created through a focused laser beam Engels and Atherton (2007); Onofrio et al. (2000); Lim et al. (2022). Accordingly, the range of the potential heights $V_{0}\in[0,5]$ and widths $\sigma=0.05$ ($0.5$) considered herein in the employed dimensionless units refer to intensities $\sim[0,240]~{}{\rm nK}$ and wavelengths $25$ $(254)~{}{\rm nm}$ respectively. Finally, evolution times $t\sim 200$ correspond to $\sim 32~{}{\rm ms}$ which are feasible in current droplet experiments Cavicchioli et al. (2025); Semeghini et al. (2018); D’Errico et al. (2019); Cheiney et al. (2018). Another interesting realization of our setup may stem from a highly particle imbalance heteronuclear three-component mixture (not realized in relevant experiments so far). Here, the droplet is again composed by two hyperfine states of 39K and the impurity by a heavy species such as 174Yb.
@@
Figure 2: Ground-state density configurations of a harmonically trapped droplet in the presence of (a)-(c) an attractive ($V_{0}<0$) and (d)-(f) a repulsive ($V_{0}>0$) potential well with respect to $V_{0}$ and different widths $\sigma$ (see legends). It is evident that the droplet shrinks in size under the influence of an attractive potential well and eventually disappears, a process that occurs faster for wider potentials. A repulsive potential leads to two droplet segments especially for larger widths. The chemical potential of the droplet and the strength of the LHY interaction are held fixed (see legends). The insets in panels (c) and (f) show the droplet density distributions when $\mu=-0.1$ and $\delta=0.7$ for the attractive and the repulsive potentials respectively. A reduced LHY strength facilitates the occurrence of the droplets for larger negative values of $V_{0}$, while splitting of the structure takes place at smaller positive $V_{0}$.
@@
To provide evidences about the many-body character of the eGPE droplet dynamics and understand its microscopic characteristics we also perform exact-diagonalization on the level of the single-particle Schrödinger equation. The respective eigenvalue problem reads $[(-1/2)\partial^{2}/\partial x^{2}+V(x)]\varphi_{n}(x)=\epsilon_{n}\varphi_{n}(x)$, where $\epsilon_{n}$, $\varphi_{n}(x)$ represent the eigenvalues and eigenstates of the system. This is numerically solved via standard diagonalization [^2] Using a spatial discretization of $dx=0.02$, guarantees numerical convergence of the eigenenergies and eigenstates of our setting. yielding $\epsilon_{n}$ and $\varphi_{n}(x)$, with the latter being, of course, orthonormal namely $\braket{\varphi_{n}}{\varphi_{m}}=\delta_{nm}$. For completeness, we remark that the kinetic energy operator is constructed with the finite difference method leading to a tridiagonal matrix representation.
@@
The first few energetically lower single-particle eigenstates of $V(x)$ with the repulsive and the attractive potential are shown in Fig. 1(c), (d). It can be seen that in the case of a double well, $V_{0}>0$, eigenenergy doublets appear with the corresponding $\varphi_{n}(x)$ being delocalized over both wells, and hence bearing similarities with Bloch states Lewenstein et al. (2012). The doublets consist of a symmetric/antisymmetric wave function pair. As expected Lewenstein et al. (2012), the energy gap decreases (increases) with $V_{0}$ for eigenstates belonging to the same (different) doublet and eventually saturates, see Fig. 1(e).
@@
On the other hand, for $V_{0}<0$ there is always a bound state (having the form of $\varphi_{0}(x)\sim\sqrt{\kappa}e^{-\kappa\left|x\right|}$, $\kappa=\left|V_{0}\right|\sigma\sqrt{2\pi}$ for $\sigma\to 0$ retrieving the bound state in a delta potential) lying within the attractive potential [Fig. 1(d)]. This is followed by delocalized states, whose extent is dictated by the harmonic oscillator length, $l_{x}=\Omega^{-1/2}=31.6$. The odd $\varphi_{n}$ eigenstates ($n=1,3,\ldots$) are not affected by the central dip since they possess a node at this location. This is manifested by their eigenenergies being close to the values in the absence of a dip, $(n+1/2)10^{-3}\hbar\omega_{\perp}$ [Fig. 1(d)]. In contrast, the even eigenstates ($n=2,4,\ldots$) are affected by the attractive potential, exhibiting a suppressed peak at the center [inset of Fig. 1(d)]. Moreover, their eigenenergies are shifted to lower values compared to the case with $V_{0}=0$. The energy gaps among successive eigenstates increase for larger $V_{0}$ and additional bound states appear as can be inferred from Fig. 1(f).

### v2

# Localization and splitting of a quantum droplet with a potential defect

@@ Abstract
We unravel the existence and nonequilibrium response of one-dimensional harmonically trapped droplet configurations in the presence of a central potential barrier or well. For fixed negative chemical potentials, it is shown that droplets fragment into two for increasing potential barrier heights, a process that occurs faster for larger widths. However, atoms from the droplet accumulate at the potential well, especially for wider ones, leading to a deformed droplet and eventually to the termination of the solution. Linearization analysis yields the underlying excitation spectrum which dictates stability and the behavior of the ensuing collective modes. Quenches in the potential height are used to demonstrate dynamical fragmentation of the droplet for potential barriers as well as self-evaporation along with droplet localization and eventual relaxation for longer evolution times in the case of potential wells. The presence of selective excitation processes emanating from quantum superposition in the induced droplet dynamics is explicated by evaluating the contribution of the participating single-particle eigenstates. Our results should be detectable by current ultracold atom experiments and may inspire engineered droplet dynamics with the aid of external potentials.
@@ I Introduction
Recent ultracold atom experiments offer access to the regime of weak attractive interactions Böttcher et al. (2020); Chomaz et al. (2023); Chen and Hung (2021). Such settings can host, among others, ultradilute liquid-type many-body droplet states Luo et al. (2021); Malomed (2021); Mistakidis et al. (2023) that are sustained due to the presence of quantum fluctuations in dipolar gases Böttcher et al. (2020); Politi et al. (2022); Chomaz et al. (2023) as well as homonuclear Cheiney et al. (2018); Semeghini et al. (2018); Cabrera et al. (2018) and heteronuclear D’Errico et al. (2019); Burchianti et al. (2020) short-range Bose mixtures. Commonly the interaction Böttcher et al. (2020) and dimension Ilg et al. (2018); Pelayo et al. (2025a) dependent Lee-Huang-Yang (LHY) first-order correction term Lee et al. (1957) to the mean-field energy suffices to capture the formation and dynamics of quantum droplets within a suitable extended Gross-Pitaevskii (eGPE) model Petrov (2015); Petrov and Astrakharchik (2016); Englezos et al. (2025). Beyond LHY effects have been, for instance, documented in Refs. Parisi et al. (2019); Mistakidis et al. (2021); Englezos et al. (2023).
@@
On the other hand, the immersion of impurity atoms into droplets can bring forth additional mechanisms including self-localization of the impurity Abdullaev and Galimzyanov (2020); Bighin et al. (2022), hybrid droplet-impurity modes Sinha et al. (2023), tunability of the dipolar bound state character Wenzel et al. (2018), and mediated impurity interactions from the droplet Pelayo et al. (2024). The impact of a heavy impurity on its host bosonic droplet remains, however, elusive. As demonstrated in repulsive Bose gases Campbell et al. (2014); Brauneis et al. (2022), the heavy impurity can be well approximated by an external potential barrier, an approach we also adopt in this work. In this context, a repulsive barrier dragged through the droplet has been used recently to dynamically generate defects like solitary waves and vortices Saqlain et al. (2023); dos Santos et al. (2025), a well-established protocol in repulsive Bose gases Engels and Atherton (2007); Hakim (1997); Rodrigues et al. (2009); Katsimiga et al. (2018) but with an amended local speed of sound. Additionally, the tunneling dynamics of droplets in a double-well potential have been discussed Abdullaev et al. (2025); Wysocki et al. (2024) as well as the transmission and reflection properties of droplets passing through a potential well have been reported Debnath et al. (2023).
@@
Accordingly, it is interesting to study the effect of the impurity onto the droplet such as to retain its bound state character in a structurally modified configuration or its destruction. Here, the computation of the underlying excitation spectrum being able to reveal the stability of the resulting states and their ensuing modes is certainly desirable. In another vein, the dynamical response of the droplet by switching-on the impurity can trigger specific excitation modes or even fragmentation. To explore these possibilities, we deploy either a potential barrier or well in a symmetric Petrov and Astrakharchik (2016) (alias single-component) one-dimensional (1D) quantum droplet featuring contact interactions. Notice that in the limiting case of a delta-like potential barrier (well) the composite system resembles an infinitely heavy impurity pinned at the trap center interacting repulsively (attractively) with its droplet host, while the impurity-droplet coupling is regulated by the value of the potential height. Similar external confinements with attractive single-particle potentials are used to emulate the motion of electrons in solids Atkins and Friedman (2011), which further promotes the broader character of our results. The static properties and nonequilibrium dynamics of this setting are modeled by the appropriate eGPE Petrov and Astrakharchik (2016).
@@
On the ground-state, we determine regions of existence of droplet states in the presence of a potential well or barrier. The former favors droplet’s spatial localization as was also argued in Ref. Debnath et al. (2023). However, here we explicate the controllability of this mechanism demonstrating, for instance, the existence of an upper bound of the height above which the droplet disappears. This threshold decreases for larger widths or greater LHY strengths, which is traced back to the appearance of additional bound states in the potential well and the increased droplet localization Petrov (2015) respectively. In contrast, a potential barrier enforces the gradual deformation and eventual fragmentation of the droplet into two, a process that occurs faster for reduced heights in the case of larger widths. We also infer the spectral (in)stability of the ensuing droplet configurations in the presence of a potential barrier or well through Bogoliubov-de Gennes (BdG) linearization analysis. The behavior of their collective modes in terms of the central potential defect characteristics is also assessed.
@@
Upon quenching a potential barrier, we showcase droplet’s fragmentation into two counterpropagating segments whose separation and velocity can be tuned by means of the potential height and width. A semiclassical model is devised to capture this response, based on energy conservation. Meanwhile, the quench excites the droplet fragments which suffer self-evaporation Ferioli et al. (2020) releasing small density portions. On the other hand, the sudden ramp-up of the potential well enforces spatial localization of the droplet which features excitation processes. The latter, due to self-evaporation, decay in the course of the evolution letting the droplet to eventually attain a quasi-steady state. Monitoring the position variance allows to track the collective motion of the droplet and its excitation frequencies whose values are in-sync with the BdG predictions. Finally, insights into the participating microscopic mechanisms and the crucial role of the quantum superposition for the droplet dynamics are unveiled through the single-particle eigenstates of the external potential. The dynamical behavior of the eigenstates inherits information about the droplet’s excitation processes.
@@
This work is organized as follows. Section II introduces our droplet setting along with its external potential and eigenstates as well as the reduced single-component eGPE framework. Section III elaborates on the characteristics of the droplet solutions and their stability in the presence of a potential well and barrier. In Section IV we analyze the emergent nonequilibrium droplet dynamics induced by suddenly switching on the potential defect which facilitates the dynamical splitting [Sec. IV.1] or localization [Sec. IV.2] of the original droplet. A decomposition of the many-body dynamics in terms of the single-particle eigenstates of the external confinement obtained through exact diagonalization is discussed in Sec. IV.3. We conclude and discuss future research extensions based on our results in Section V.
@@ II Droplet setting and external potential
[…]
The mixture is confined by a harmonic oscillator with frequency $\omega_{x}\ll\omega_{\perp}$, where $\omega_{x}$ ($\omega_{\perp}$) denotes the trapping frequency in the elongated (transverse) direction. This condition together with $|\mu|/\hbar\ll\omega_{\perp}$ (with $\mu$ denoting the droplet’s chemical potential) ensure the 1D character of the system. Namely, due to tight confinement the ensuing excitation energies along the perpendicular $y$, $z$ directions are much larger than the elongated excitation energies. Hence, the atoms are kinematically constrained along the $x$ direction. Accordingly, the 1D droplet, in the weakly interacting regime, is adequately described by the eGPE Petrov and Astrakharchik (2016); Englezos et al. (2025)
@@
experiencing the competition of repulsive cubic mean-field interactions and attractive quantum fluctuations scaling as $\Psi^{2}$. The latter are responsible for droplet formation Petrov and Astrakharchik (2016); Luo et al. (2021). Here, $\Psi(x,t)$ refers to the 1D droplet wave function and $\hbar$ is the reduced Planck constant. The participating interaction strengths are experimentally tunable through manipulation of the three-dimensional scattering lengths that can be achieved by means of available magnetic fields using Fano-Feshbach resonances Semeghini et al. (2018); Cheiney et al. (2018) or via confinement induced resonances Olshanii (1998); Bergeman et al. (2003) which can be attained by adjusting the transverse confinement.
Figure 1: Schematic illustration of the external potential described by Eq. (2) corresponding to (a) a double-well with $V_{0}>0$ and (b) a harmonic trap with a central dip when $V_{0}<0$. The single-particle energy levels are marked by $\epsilon_{n}$, with $n=0,1,2,\dots$, obtained through exact diagonalization. The shaded areas in panel (a) designate the energy doublets. The respective eigenstates are provided in panels (c) [(d)] for $V_{0}>0$ $V_{0}<0$]. In both cases, the employed potential characteristics correspond to $V_{0}=\pm 1$ and $\sigma=0.5$. (e), (f) The dependence of the underlying energy gaps among consecutive energy levels, $\Delta\epsilon_{n}=\epsilon_{n}-\epsilon_{n-1}$, with $n=1,2,\dots$ as a function of the potential depth $V_{0}$ (see legend). Inset in panel (d) shows a magnification at the center of $\left|\varphi_{4}\right|^{2}$ to clearly visualize the existent “kink” of even eigenstates. In the double-well, $\Delta\epsilon_{n}$ among doublets increase for larger $V_{0}$, while in the potential well case additional bound states (BS) appear for increasing $V_{0}$ (shaded regions).
The external potential consists of a loose harmonic trap [^1] Note that the presence of the harmonic trap is not necessary for the results to be presented below. Similar phenomena occur also with a box confinement in the $x$-direction. with frequency $\Omega=\omega_{x}/\omega_{\perp}=0.001$ and a superimposed central defect
@@
Here, $\sigma$ denotes the spatial width of the central potential defect, and $V_{0}$ refers to the strength of the potential well ($V_{0}<0$) or barrier ($V_{0}>0$). Apparently, the central potential defect reduces to a delta-like one for $\sigma\to 0$ and $V_{0}\to\pm\infty$, while $V_{0}\sigma$ is kept fixed. Specifically, for $V_{0}<0$, the overall confinement corresponds to a harmonic trap with a central dip, see Fig. 1(b), facilitating the accumulation of atoms from the droplet at its vicinity since it hosts additional bound states depending on $V_{0}$ (see also the discussion below). This contribution acts against flat-top configurations and its increase may gradually destruct the droplet. The potential well hosts a larger number of bound states for larger $V_{0}$ [Fig. 1(f)], while their energy gap is reduced for smaller $\sigma$. On the other hand, for $V_{0}>0$ a 1D double-well potential [Fig. 1(a)] is realized which favors droplet fragmentation into two segments. In this case, a larger $V_{0}$ results in the increase of the energy gap among consecutive single-particle eigenstate doublets, see also Fig. 1(e).
@@
where $\lambda=\frac{\delta g}{2}\sqrt{(\frac{m}{\omega_{\perp}\hbar^{3}})}$, and $\delta=\frac{1}{\pi}(\frac{mg^{2}}{\omega_{\perp}\hbar^{3}})^{3/4}$ represent respectively the strength of the mean-field and the LHY quantum correction. Also, $V(x)=\frac{1}{2}\Omega^{2}x^{2}+\frac{V_{0}e^{-x^{2}/{2\sigma^{2}}}}{\hbar\omega_{\perp}}$. As such, all quantities to be presented below are in dimensionless units. Since we aim to examine the non-trivial impact of the potential defect on the droplet, barrier heights satisfying $\left|V_{0}\right|\gg\left|\mu\right|$ (with $\mu$ denoting the droplet’s chemical potential) are considered. Also, the healing length of the droplet, $\xi=\sqrt{\lambda}/\delta=1$, sets a characteristic length scale. Accordingly, the central potential widths examined obey $\sigma<\xi$; otherwise droplet destruction occurs already for very small $V_{0}$ values.
@@
Figure 2: Ground-state density configurations of a harmonically trapped droplet in the presence of (a)-(c) a potential well ($V_{0}<0$) and (d)-(f) a barrier ($V_{0}>0$) with respect to $V_{0}$ and different widths $\sigma$ (see legends). It is evident that the droplet shrinks in size under the influence of a potential well and eventually disappears, a process that occurs faster for wider potentials. A potential barrier leads to two droplet segments forming faster especially for larger widths. The chemical potential of the droplet and the strength of the LHY interaction are held fixed (see legends). The insets in panels (c) and (f) show the droplet density distributions when $\mu=-0.1$ and $\delta=0.7$ for the potential well and barrier respectively. A reduced LHY strength facilitates the occurrence of the droplets for larger negative values of $V_{0}$, while splitting of the structure takes place at smaller positive $V_{0}$.
Typical trap ratios $\Omega=0.001$ can be achieved through tight transversal confinement $\omega_{\perp}/(2\pi)=1$~kHz which is in a similar range to the 1D experiments of Refs. Görlitz et al. (2001); Moritz et al. (2003); Meinert et al. (2017); Romero-Ros et al. (2024) and frequency $\omega_{x}/(2\pi)=1$~Hz along the elongated axis. The central potential defect is created through a focused laser beam Engels and Atherton (2007); Onofrio et al. (2000); Lim et al. (2022). Accordingly, the range of the potential heights $V_{0}\in[0,5]$ and widths $\sigma=0.05$ ($0.5$) considered herein in the employed dimensionless units refer to intensities $\sim[0,240]~{\rm nK}$ and wavelengths 2525 $(254)~{\rm nm}$ respectively. Finally, evolution times $t\sim 200$ correspond to $\sim 32~{\rm ms}$ which are feasible in current droplet experiments Cavicchioli et al. (2025); Semeghini et al. (2018); D’Errico et al. (2019); Cheiney et al. (2018). Another interesting realization of our setup may stem from a highly particle imbalance heteronuclear three-component mixture (not realized in relevant experiments so far). Here, the droplet is again composed by two hyperfine states of 39K and the impurity by a heavy species such as 174Yb.
@@
To provide evidences about the many-body character of the eGPE droplet dynamics and understand its microscopic characteristics we also perform exact-diagonalization on the level of the single-particle Schrödinger equation. The respective eigenvalue problem reads $(-1/2)\partial^{2}/\partial x^{2}+V(x)]\varphi_{n}(x)=\epsilon_{n}\varphi_{n}(x)$, where $\epsilon_{n}$, $\varphi_{n}(x)$ represent the eigenvalues and eigenstates of the system. This is numerically solved via standard diagonalization [^2] Using a spatial discretization of $dx=0.02$, guarantees numerical convergence of the eigenenergies and eigenstates of our setting. yielding $\epsilon_{n}$ and $\varphi_{n}(x)$, with the latter being, of course, orthonormal namely $\braket{\varphi_{n}|\varphi_{m}}=\delta_{nm}$. For completeness, we remark that the kinetic energy operator is constructed with the finite difference method leading to a tridiagonal matrix representation.
@@
The first few energetically lower single-particle eigenstates of $V(x)$ with the potential barrier and well are shown in Fig. 1(c), (d). It can be seen that in the case of a double well, $V_{0}>0$, eigenenergy doublets appear with the corresponding $\varphi_{n}(x)$ being delocalized over both wells, and hence bearing similarities with Bloch states Lewenstein et al. (2012). The doublets consist of a symmetric/antisymmetric wave function pair. As expected Lewenstein et al. (2012), the energy gap decreases (increases) with $V_{0}$ for eigenstates belonging to the same (different) doublet and eventually saturates, see Fig. 1(e).
@@
On the other hand, for $V_{0}<0$ there is always a bound state lying within the potential well [Fig. 1(d)]. This bound state has the form $\varphi_{0}(x)\sim\sqrt{\kappa}e^{-\kappa\left|x\right|}$, with $\kappa=\left|V_{0}\right|\sigma\sqrt{2\pi}$ which in the limiting case of ($\sigma\to 0$, $V_{0}\to\infty$) converges to the relevant bound state in a delta potential. Higher-excited states are delocalized, with their extent being dictated by the harmonic oscillator length, $l_{x}=\Omega^{-1/2}=31.6$. The odd $\varphi_{n}$ eigenstates ($n=1,3,\ldots$) are not affected by the central dip since they possess a node at this location. This is manifested by their eigenenergies being close to the values in the absence of a dip, $(n+1/2)10^{-3}\hbar\omega_{\perp}$ [Fig. 1(d)]. In contrast, the even eigenstates ($n=2,4,\ldots$) are affected by the potential well, exhibiting a suppressed peak at the center [inset of Fig. 1(d)]. Moreover, their eigenenergies are shifted to lower values compared to the case with $V_{0}=0$. The energy gaps among successive eigenstates increase for larger $V_{0}$ and additional bound states appear as can be inferred from Fig. 1(f).